Which equation represents the line that passes through (0,0) and is parallel to the line passing through (2,3) and (6,1)?
a) y= 1/2x
b) y= -1/2x
c) y= -2x
d) y= 2x

Answer by stanbon(75887)

Which equation represents the line that passes through (0,0) and is parallel to the line passing through (2,3) and (6,1)?
-------------
Since f(0)= 0, the intercept is 0.
----------
Those points yield a slope of (1-3)/(6-2) = -2/4 = -1/2
--------------
Ans: b
=============
Cheers,
Stan H.
a) y= 1/2x
b) y= -1/2x
c) y= -2x
d) y= 2x